my question is as follows pls help me i cant solve it using distributive property
1427 x 99 - (- 1427)
 

Grade:10

1 Answers

Arun
4 years ago
Dear student
 
1427 * 99 – ( – 1427) = 1427* 99 + 1427 = 1427 (99+1) = 1427* 100 = 142700
 
Hope it helps
 
Regards
Arun